OpenStudy (anonymous):

Identify the discontinuities, graph, and find the zeros of the rational function f(x) = (4x^2-36x)/(x-9)

OpenStudy (anonymous):

\[\frac{4x^2-36x}{x-9}=\frac{4x(x-9)}{x-9}\]

OpenStudy (anonymous):

when is this problem undefined

OpenStudy (anonymous):

got that part i just dont know what the point of discontinuity is... all i know is that its going to bo (9,y)

OpenStudy (anonymous):

undefined?

OpenStudy (anonymous):

when is there an asymptote

OpenStudy (anonymous):

undefined is when you have a denominator that equals 0 or when a function is undefined such as ln(-1)

OpenStudy (anonymous):

9? x cant equal 9?

OpenStudy (anonymous):

yes so 9 is a discontinuity

OpenStudy (anonymous):

because at that point it breaks from the continuous line. it's an asymptote now to find out if it's a removable discontinuity, simplify the problem... if you cancel out the (x-9)'s

OpenStudy (anonymous):

4x. but i thought the discontinuity was a specific point....

OpenStudy (anonymous):

it is so at x=9 you have the function 4x which is continuous for all real so now you have a piecewise function as such